# Memorization to Generalization: Emergence of Diffusion Models from Associative Memory

### v2

@@ 3 Memorization-to-Generalization in 2D
Notice, the dependence on the polar angle $\phi$ in Equation 12 completely disappears from the final result. The local minima of the resulting energy function form a continuous manifold, corresponding to $R=1$. Data samples from the model occupy the vicinity of that manifold. This behavior describes the fully generalized phase. In the limit of large $\beta$ the energy landscape is described by a parabola centered around $R=1$.
For a single training data point, this simple model exhibits memorization. With a few training data points, the model can either be in the memorization or spurious phase, depending on the location of those points and the value of $\beta$. For a very large number of training data points, the model is in the generalization phase. All effects perfectly align with the cartoon illustration in Figure 1.
Overall, this calculation entirely ignores the neural network, which is used to model the score and approximate the empirical energy. There exists a logical possibility that the introduction of the neural network would entirely change the progression of distinct behaviors: memorization, spurious, generalization. To test this, we trained a set of diffusion models in this precise setting. Each model is trained on a dataset of size $K$. The resulting energy landscapes are shown in Figure 3. For small number of data points, e.g., $K=2$, the diffusion model exhibits memorization. The local minima of the energy correspond to the training data points. Importantly, at $K=9$, we are able to observe the first signs of spurious states. At this stage, the model begins to learn emergent (different from the training data) local minima of the energy. Subsequent increase of the size of the training set leads to fully generalized behavior, which is illustrated for $K=1000$ data points. At that stage, all of the samples from the model live in close proximity of the exact data manifold. The right panel shows the analytical expression for the energy landscape, defined by Equation 12. Thus, the conventional diffusion modeling pipeline agrees very well with the theoretical prediction of the empirical energy. Please refer to Appendix B for full details on the toy model experiment.
4 Detection Metrics
Figure 4: Different sample types across the memorization-to-generalization transition for CIFAR10 [63]. The grey histogram shows the distances between synthetic samples and their nearest neighbors from the synthetic set $\mathsf{S}^{\prime}$. The threshold $\delta_{s}$ defines a boundary between the two peaks. The olive histogram depicts the distances from the synthetic samples to their closest neighbor from the training set $\mathsf{S}$, with threshold $\delta_{m}$ separating the two peaks. Memorized samples are located in the left peak of the olive histogram, below $\delta_{m}$. In contrast, generalized and spurious samples appear to the right of $\delta_{m}$ in the olive histogram. Examples of the generated samples forming each of the four peaks of the histograms are shown in the inset frames. For each generated sample, top-4 nearest neighbors from the training set are shown in the top row, and those from the synthetic set are shown in the bottom row. Training set size $K=7310$ was used in this figure (peak of the frequency of spurious states), but the discussed phenomena are general and largely independent of this specific value. The fraction of the memorized, spurious, and generalized samples in the pool of all generated samples is shown in the bottom left panel as a function of the training set size. The inset shows amplified spurious fraction (green curve).
The main goal of our work is to establish the existence of spurious states in diffusion models trained on natural high-dimensional datasets, not just toy model problems. With this goal in mind we design detection metrics that can classify any generated sample into one of the three categories: memorized, spurious, and generalized samples. Fundamentally, these metrics defined below rely on two datasets: $\mathsf{S}$ — the training dataset used to train the diffusion model, and $\mathsf{S}^{\prime}$ — the synthetic dataset generated from the fully trained diffusion model. The size of the synthetic set $\mathsf{S}^{\prime}$ is assumed to be much bigger than the training set size.
The core intuition comes from the typical energy landscape around these three distinct kinds of generated samples, see Figure 1. Memorized samples are elements of the synthetic set $\mathsf{S}^{\prime}$, which have duplicates in the training set $\mathsf{S}$. Spurious samples are elements of the synthetic set $\mathsf{S}^{\prime}$, which have duplicates in the synthetic set $\mathsf{S}^{\prime}$, but do not have duplicates in $\mathsf{S}$. In other words, the same spurious sample appears in the synthetic set more than once. Finally, generalized samples are unique elements of $\mathsf{S}^{\prime}$, which do not have duplicates in either the synthetic set $\mathsf{S}^{\prime}$ nor the training set $\mathsf{S}$.
With this intuition in mind we examine two histograms, see Figure 4. First, for every element of the synthetic set $\mathsf{S}^{\prime}$ the distance to the closest nearest neighbor from the training set $\mathsf{S}$ is shown by the olive histogram. Second, for every element of the synthetic set $\mathsf{S^{\prime}}$ the distance to the closest nearest neighbor in the synthetic set $\mathsf{S}^{\prime}$ is shown by the gray histogram. Both histograms exhibit a clear bimodal shape, indicating that at least two groups of samples are present in the synthetic set. This bimodal structure makes it possible to select the memorized $\delta_{m}$ and spurious $\delta_{s}$ thresholds to separate the samples forming the two peaks in each histogram. The left peak of the olive histogram is composed of synthetic samples that are at a small distance away from the training set $\mathsf{S}$, which can be considered as memorized samples. The left peak of the gray histogram contains synthetic samples that have duplicates in the synthetic set $\mathsf{S}^{\prime}$, indicating they can be memorized if the same sample appears in the training set, or otherwise, spurious. The right peak of the gray histogram contains generalized samples, and the right peak of the olive histogram contains generalized and spurious samples. These observations make it possible to formally define the following detection metrics.
Metric 4.1 (Memorization Detection).
Following [45], we define the memorization detection metric $\mathcal{M}$. Given a target pattern $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ and its first nearest neighbor ${\mathbf{x}}_{1}$ extracted from the training set $\mathsf{S}$ according to a distance measure $d(\cdot,\cdot)$, where a small distance corresponds to high similarity, metric $\mathcal{M}$ detects sample $\hat{{\mathbf{x}}}$ as memorized if $d(\hat{{\mathbf{x}}},{\mathbf{x}}_{1})$ is small.
$\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})=\mathbb{I}\bigg{(}d(\hat{{\mathbf{x}}},{\mathbf{x}}_{1})\leq\delta_{m}\bigg{)}$ (13)
where $\mathbb{I}$ represents the indicator function and $\delta_{m}\in\mathbb{R}$ is a threshold defined by the distance histograms, see Figure 4.
Metric 4.2 (Spurious Detection).
Given a synthetic sample $\hat{{\mathbf{x}}}$ and its first nearest neighbor from the synthetic set ${\mathbf{x}}^{\prime}_{1}\in\mathsf{S}^{\prime}$, the spurious pattern detection metric $\mathcal{S}$ identifies instances where the model generates outputs that do not belong to the training set but have high similarity with samples from the synthetic set (left peak of the gray histogram in Figure 4 after the exclusion of memorized samples).
$\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\mathbb{I}\bigg{(}d(\hat{{\mathbf{x}}},{\mathbf{x}}^{\prime}_{1})\leq\delta_{s}\bigg{)}\land\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})$ (14)
where $\delta_{s}\in\mathbb{R}$ is a threshold value chosen based on the gray distance histogram, see Figure 4.
Metric 4.3 (Generalization Detection).
In contrast, the sample $\hat{{\mathbf{x}}}$ is generalized if it is neither memorized nor spurious.
$\mathcal{G}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})\land\neg\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})$ (15)
5 Memorization - Spurious - Generalization Transition
Figure 5: Fractions of memorized, spurious, and generalized samples in synthetic sets across training sizes and datasets. As the training data size $K$ increases, memorization decreases and the fraction of generalized samples steadily increases (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row).
Using our detection metrics, we computed the fractions of memorized, spurious, and generalized samples for various sizes of training data for different datasets, see Figure 5. These datasets include MNIST [64], FASHION-MNIST [65], CIFAR10 [63], and LSUN-CHURCH [66] scaled down to $64\times 64$ resolution using center-crop and down-scale. For each dataset, we trained DDPM-based diffusion models [23] for $M=38$ different training set sizes. For each model $\alpha=1,...,M$, trained on the training set $\mathsf{S}_{\alpha}$, a synthetic set $\mathsf{S}^{\prime}_{\alpha}$ was generated. To account for duplication, we ensured that each $\mathsf{S}^{\prime}_{\alpha}$ is four times the size of its corresponding $\mathsf{S}_{\alpha}$, and checked that further increase of the synthetic set size does not qualitatively change our results, see Figure 11 in Appendix D. Each sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}_{\alpha}$ is classified as either memorized, spurious, or generalized using the above metrics (13), (14), and (15). The fractions of these three pattern types with respect to the size of $\mathsf{S}^{\prime}_{\alpha}$ are plotted in Figure 5. For each dataset, our smallest model was trained on the training set of $|\mathsf{S}_{1}|=2$ data points while the largest model was trained on the entire original training set $\mathsf{S}_{M}=\mathsf{S}$. Please refer to Appendix D for additional details on training, selection of data sizes, and detection metrics’ hyperparameters for each dataset.
The results in Figure 5 clearly demonstrate the transition from memorization to generalization as the dataset size increases. Meanwhile, the collected samples also show distinct characteristics in each of the considered phases, see Figure 2. Specifically, in the small data regime, we see in Figure 5 that the diffusion model predominantly replicates the training data (see the memorized panel of Figure 2, or additional samples in Figure 15 of the Appendix). As the data size surpasses the memorization capacity, e.g., at $K=3931$ for CIFAR10 in Figure 5, we observe a critical transition where the memorization fraction declines and spurious patterns become prominent. Such patterns exhibit strong duplication in the synthetic set (see the spurious panel of Figure 2 and additional samples in Figure 16 of the Appendix). The emergence of such patterns aligns with the onset of generalization, as the model moves away from strictly reproducing the training set and starts generating novel combinations of learned features. Subsequent increase of the training set size leads to the decrease of the fractions of memorized and spurious samples, signaling the transition to the full generalization regime. At this stage, the model completely loses its replication ability and no duplicate samples are detected in either training or synthetic sets (see the generalized panel of Figure 2 and additional samples in Figure 17 of the Appendix). Overall, this particular trend is also observed as the size of the diffusion model’s neural network is varied, see the results in Figure 9 of Appendix D.
6 The Basin of Attraction Across the Transition
Figure 6: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in CIFAR10 [63], computed with up to 512 samples per type. When fewer than 512 samples are available, all points are used. The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models. Shaded regions indicate standard deviation of the log-volume values.
Diffusion models provide us with an opportunity to define the notion of basins of attraction around generated samples to further strengthen their connection to associative memories. Specifically, as the training data size $K$ varies, the sizes of those basins for memorized, spurious, and generalized samples can exhibit interesting behaviors. This section studies these phenomena empirically.
Inspired by [38, 67, 54], we perturb a pattern ${\mathbf{x}}_{0}$ through the forward diffusion process for $T$ steps, and attempt to reverse it back through the reverse diffusion process. The maximum or critical time $t_{c}$, where the reverse process can still recover the pattern ${\mathbf{x}}_{0}$ from its perturbation ${\mathbf{x}}_{t_{c}}$ within an error margin $\delta\in\mathbb{R}$, defines the basin of attraction for pattern ${\mathbf{x}}_{0}$. The radius of this basin $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})=\lVert{\mathbf{x}}_{0}-{\mathbf{x}}_{t_{c}}\rVert$ is the Euclidean norm of the difference between ${\mathbf{x}}_{0}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$. This radius can be converted to volume assuming that the basin of attraction is a sphere of radius $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})$. Please also refer to Appendix G for more details on the experimentation and computation of $t_{c}$ and the volume.
The results for log-volume versus training size are shown in Figure 6. The basins of attraction for training data decrease as $K$ increases. This trend is expected, since packing larger amount of training data in a given volume of the configuration space results in a smaller volume per data point, similarly to the ideal gas in physics. Memorized samples follow a similar trend. These results suggest that as $K$ grows, the diffusion model becomes less effective at denoising the perturbed data points, and instead gets much better at synthesizing. The most significant finding is that each of the three sample types exhibits a distinct log-volume trend: memorized samples have the largest log-volume, generalized samples have the smallest log-volume (this is expected as they should be unlikely recoverable by definition, resulting in close to zero volume), and spurious samples fall in between the two. This finding suggests a clear separation in how different sample types behave as the model’s training data size progresses. Please refer to Figs. (18) and (19) in the Appendix for the results of average log-volume across the transition for the other datasets, which exhibit the same behavior as Figure 6.
To further illustrate the shrinkage of basins of attraction, we visualize the degradation in the denoising process as the training data size $K$ increases, as shown in Figure 7. For each training image, we apply the forward diffusion process using a timestep $t\approx t_{c}$, chosen near the end of the memorization phase (see Figure 20 in the Appendix). The perturbed image is then denoised using the deterministic DDIM sampler [24], starting from that same timestep $t$, across diffusion models trained with varying $K$. When $K$ is small, the diffusion model behaves like a DenseAM, effectively denoising and recovering the original image. However, as $K$ increases, this retrieval ability weakens, and the model shifts toward generating novel samples. In the intermediate regime, before full generalization, the outputs of the denoising process visually resemble spurious patterns: incomplete or noise-like novel generations which indicate the onset of generalization. Please refer to Appendix G for more details on this visual demonstration.
Figure 7: An illustration of the shrinkage of basins of attraction as the training data size increases. Two CIFAR10 [63] and two LSUN-CHURCH [66] training images are independently perturbed using different timesteps $t\approx t_{c}$ (shown on the right), each chosen near the end of the memorization phase (see Figure 20 in the Appendix), and then denoised using the deterministic DDIM sampler [24]. At small training data sizes, the model accurately recovers the original image. As the training data size increases, denoising performance deteriorates: the model shifts from retrieval of data points to generating novel outputs. In intermediate regimes, the outputs resemble spurious patterns: partial and noise-like generations indicative of this transition.
7 The Relative Energy Gap Across the Transition
Figure 8: Average relative energy gap of each sample type and the training samples in CIFAR10 [63] as training data size $K$ grows. The gaps are measured relative to training samples, whose energy is set to zero (energy is defined up to an additive constant). Shaded regions show standard deviation of the gap values.
Lastly, we study how the energy of the three sample types changes as the training data size increases. Following the efficient relative energy computation from [38], we measure the relative energy gap of the three sample types with respect to that of the training data samples. Please refer to Appendix H for full details on the experiment and computation of the relative energy gap for different sample types.
The results are recorded in Figure 8. For small training data sizes, the energy of the memorized samples is low and of the generalized samples is high. As $K$ increases, the energy of the memorized samples increases, while the energy of the generalized samples decreases. The energy of the spurious states is the lowest among these three types of sample at the memorization-generalization transition. Please refer to Figure 23 in the Appendix for the results of the average relative energy gap across the transition for the other datasets, which show a similar behavior as Figure 8.
8 Conclusion
In this work, we established a novel connection between diffusion models and DenseAMs. Specifically, we identified the emergence of creativity (generalization) in diffusion models with the failure of successful memory recalls, via an energy-based theoretical description of diffusion models. Using this theory, we identified a novel phase – spurious states – previously overlooked in the memorization-generalization literature. Finally, we established that diffusion models trained on natural datasets follow the same transition (memorization to spurious phase to generalization) as predicted by our theory. The peak frequency of spurious states among generated samples marks the onset of generalization as the training set size increases. The volume measurements and relative energy gap analysis of different sample types, across multiple datasets consistently support this three distinct phases of computation, providing a comprehensive understanding of how diffusion models evolve from memorization to generalization as the amount of training data increases.
Aknowledgments
M.N. acknowledges the support of PNRR MUR project PE0000013-FAIR.
B.P. acknowledges the support of IBM through the IBM-Rensselaer Future of Computing Research Collaboration (FCRC).
Appendix A Additional Details on the Energy Formulation of the 2D Toy Example
@@
In order to obtain Equation 11, it is easiest to introduce polar coordinates for both the state vector ${\mathbf{x}}$ and the training data ${\mathbf{y}}$:
@@
[…]
@@
Appendix B Experimental Details of the 2D Toy Example
This section describes the two-dimensional toy example used in Section 3, including details about the neural network architecture, training, generation, and computation of the model’s energy. All of the corresponding code will be made publicly available.
Dataset. To construct the dataset, we uniformly sample points on the unit circle. Concretely, we take ${\mathbf{y}}\sim p({\mathbf{y}})=\frac{1}{2\pi}\delta(r-1)$, where $r=\sqrt{y_{1}^{2}+y_{2}^{2}}$ is the radius and $\delta(r-1)$ is the delta function ensures that all probability mass lies on radius $r=1$. The factor $\frac{1}{2\pi}$ guarantees that the angles are sampled uniformly, relying on the identity $\delta(\sqrt{y_{1}^{2}+y_{2}^{2}}-1)=2\delta(y_{1}^{2}+y_{2}^{2}-1)$ in $p({\mathbf{y}})=\frac{1}{\pi}\delta(y^{2}_{1}+y_{2}^{2}-1)$. In practice, we sample polar coordinates $(r,\varphi)$ with $r=1$ and uniformly sample the angular coordinate $\varphi$ from $[0,2\pi]$. We then convert $(r,\varphi)$ to Cartesian coordinates ($y_{1},y_{2}$) using $y_{1}=r\cos\varphi$ and $y_{2}=r\sin\varphi$.
We first created a data set of $60000$ points using a fixed random seed, then built smaller subsets (e.g., $K\in\{2,4,9,1000,10000\}$) by progressively adding distinct samples without replacement, ensuring that each subset is a strict subset of the next. This approach allows us to systematically examine how the model behaves, as we vary the size of the training set.
Exact Score. Following Equation 12 specified in Section 3, energy can be written in closed form, whose local minima form the continuous data manifold at $R=1$ (throughout this paper we use $r$ for the polar radius of the training data and $R$ for the polar radius of the generated samples). Furthermore, in the case of many data points or $K\rightarrow\infty$, we have $p({\mathbf{y}})=\frac{1}{\pi}\delta(y_{1}^{2}+y_{2}^{2}-1)$. Since $p(R)\propto\exp^{-E^{\text{AM}}(R)}$, the score can be described as
$\begin{split}-\nabla_{R}\log p(R)&=\nabla_{R}E^{\text{AM}}(R)\\ &=2R-\frac{2\beta I_{1}(2\beta R)}{\beta I_{0}(2\beta R)}\\ &=2R-2\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}\end{split}$ (16)
where $E^{\text{AM}}(R)$ is Equation 12, and $I_{1}$ and $I_{0}$ are the first-order and zero-order modified Bessel functions of the first kind, respectively. With further expansion, the score becomes
$\nabla_{R}\log p(R)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\qquad\Rightarrow\qquad\nabla_{x}\log p(x)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\frac{x}{R}$ (17)
We implemented Equation 12 on a $20\times 20$ grid to obtain the exact energy and score of the model, overlaying 1000 samples from the training set. For visualization purposes, we set $\beta=20$, ensuring a well-defined energy landscape. Additionally, we normalized the energy by subtracting its minimum value to ensure that the lowest energy point remains at zero.
VE-SDE. To align with the theoretical framework, we employ a variance exploding (VE) SDE of the form
@@ Appendix B Experimental Details of the 2D Toy Example
where the diffusion coefficient $g(t)=\sigma$. Assuming $t\in(0,1]$, the variance of the diffusion kernel is given by $\int_{0}^{t}g^{2}(s)\,\mathrm{d}s=\sigma^{2}t$, which allows us to construct the corresponding Gaussian kernel
$p({\mathbf{x}}_{t}|{\mathbf{y}})=\mathcal{N}({\mathbf{x}}_{t};{\mathbf{y}},\sigma^{2}t\mathbf{I})$
@@
which shares the same marginal distributions as the SDE in Equation 19 and is useful for likelihood estimation.
Model and Training Details. For each training size $K$ in Figure 3, we use a U-Net inspired MLP-based network to estimate the score $s_{\theta}({\mathbf{x}}_{t},t)$, without using any down- or up- sampling blocks. The general architecture includes (1) a Fourier random feature (FRF) timestep embedding layer and a linear layer which projects the concatenation of FRF timestep embedding and input into a latent dimension of 256; (2) an encoder which consists of four non-convolutional residual blocks (using the same latent dimension) with Swish [69] activation in between; (3) a decoder which consists of the same number of residual blocks; and (4) a linear block which projects the latent variable back to the 2D space.
The network was trained in continuous time (with 1000 discretized steps) with $\sigma=1$ using the objective function from [23, 25]. Optimization is performed using the Adam [70] optimizer with a learning rate $lr=10^{-4}$. The batch size is set to $\textbf{min}(K,500)$. All models are trained for $800000$ iterations with the maximum batch size of $500$. Meanwhile, we use Euler-Maruyama discretization method [71] to solve Equation 19 of the reverse SDE for the generation of samples.
Empirical distribution. We represent the data as an empirical distribution
@@
where $N$ is the dimensionality of the data point ${\mathbf{y}}$. Since ${\mathbf{x}}_{t}$ is drawn from the forward process distribution $p({\mathbf{x}}_{t}|{\mathbf{y}})$ that is conditioned on the data point ${\mathbf{y}}$, which can be expressed as a Gaussian kernel $\mathcal{N}({\mathbf{x}}_{t};\mathbf{y},\sigma^{2}t\mathbf{I})$. We can obtain the distribution $p(\mathbf{x}_{t},t)$ as
$\begin{split}p(\mathbf{x}_{t},t)=\int p(\mathbf{x}_{t}\mid\mathbf{y})\,p(\mathbf{y})\,\mathrm{d}\mathbf{y}&=\frac{1}{K}\sum_{\mu=1}^{K}\int\mathcal{N}\!\bigl{(}\mathbf{x}_{t};\mathbf{y},\sigma^{2}t\mathbf{I}\bigr{)}\,\delta^{(N)}\!\bigl{(}\mathbf{y}-\bm{\xi}^{\mu}\bigr{)}\,\mathrm{d}\mathbf{y}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\mathcal{N}\!\bigl{(}\mathbf{x}_{t};\,\bm{\xi}^{\mu},\,\sigma^{2}t\,\mathbf{I}\bigr{)}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\frac{1}{(2\pi\,\sigma^{2}t)^{\frac{N}{2}}}\exp\!\Bigl{(}-\,\frac{\|\mathbf{x}_{t}-\bm{\xi}^{\mu}\|_{2}^{2}}{2\,\sigma^{2}t}\Bigr{)}\end{split}$ (22)
which is Equation 4, when the variance of the Gaussian kernel is $\sigma^{2}t$.
Likelihood Computation. Following [25] we can compute the log-likelihood $\log p_{\theta}({\mathbf{x}}_{0})$ given by a diffusion model, with the instantaneous change of variable formula [72], where ${\mathbf{x}}_{0}$ are the generated samples from the model. Replacing Equation 19 into the log-likelihood equation,
$\log p_{0}({\mathbf{x}}_{0};\mathbf{\theta})=\log p_{T}({\mathbf{x}}_{T})+\int_{0}^{T}\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t,\qquad\text{ with }\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)=-\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ (23)
The function $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ came from the above Probability Flow ODE (Equation 20), and $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ denotes the divergence of the function (or the trace of its Jacobian).
To estimate the likelihood of the model, we integrate $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ from a small time ${\epsilon}=10^{-5}$ to $T=1$ using a numerical integrator and add the prior logarithmic likelihood to it following Equation 23. The divergence term $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ is computed using the Laplacian, which is computationally feasible for this toy example, instead of using the Hutchinson trace estimator [73] done in [25]. We use the RK45 method [74] implemented in the ODE solver scipy.integrate.solve_ivp from Scipy [75] to solve the above integral.
Energy Computation. To compute the energy, we use Equation 5, which can be also computed using Equation 23. From the Boltzmann distribution $p({\mathbf{x}})=\frac{1}{Z}\exp[-\beta E(\mathbf{x})\bigl{]}$, we obtained $\log p({\mathbf{x}})\propto-\beta E(\mathbf{x})$. In our case the inverse temperature is $\beta=\frac{1}{2\sigma^{2}t}$. The equation for our VE-SDE thus becomes

### v3

@@ A.1  Overview
where $R$ is the radius of the unit circle and $\phi$ is the polar angle. The dependence on the polar angle $\phi$ in Eq. (15) completely disappears from the final result, since the local minima of the resulting energy function form a continuous manifold, described by a parabola centered around $R=1$ as $\beta\rightarrow\infty$. This behavior describes the fully generalized phase.
@@
In order to obtain Eq. (14), it is easiest to introduce polar coordinates for both the state vector $\mathbf{x}$} and the training data $\mathbf{y}$}:
@@
$p({\mathbf{x}})\sim\int\limits_{0}^{2\pi}d\varphi\int\limits_{0}^{\infty}rdr\frac{1}{\pi}\delta(r^{2}-1)e^{-\beta[R^{2}+r^{2}-2Rr\cos(\varphi-\phi)]}=e^{-\beta(R^{2}+1)}I_{0}(2\beta R)$
@@
A.2  Experimental Details
Dataset. To construct the dataset, we uniformly sampled points on the unit circle. Concretely, we had ${\mathbf{y}}\sim p({\mathbf{y}})=\frac{1}{2\pi}\delta(r-1)$, where $r=\sqrt{y_{1}^{2}+y_{2}^{2}}$ is the radius and $\delta(r-1)$ is the delta function ensures that all probability mass lies on radius $r=1$. The factor $\frac{1}{2\pi}$ guarantees that the angles are sampled uniformly, relying on the identity:
$\delta(\sqrt{y_{1}^{2}+y_{2}^{2}}-1)=2\delta(y_{1}^{2}+y_{2}^{2}-1),$ (16)
where $p({\mathbf{y}})=\frac{1}{\pi}\delta(y^{2}_{1}+y_{2}^{2}-1)$. In practice, we sampled polar coordinates $(r,\varphi)$ with $r=1$ and uniformly sampled the angular coordinate $\varphi\in[0,2\pi]$. We then converted $(r,\varphi)$ to Cartesian coordinates ($y_{1},y_{2}$) using $y_{1}=r\cos\varphi$ and $y_{2}=r\sin\varphi$.
Then, We created a data set of $60000$ points using a fixed random seed, then built smaller subsets by progressively adding distinct samples without replacement, ensuring that each subset is a strict subset of the next. This approach allowed us to systematically examine how the model behaves as we vary the size of the training set.
Exact Score. Following Eq. (15), the energy can be written in closed form, whose local minima form the continuous data manifold at R=1R=1111Please note throughout this paper we use $r$ for the polar radius of the training data and $R$ for the polar radius of the generated samples.. Furthermore, in the case of many data points or $K\rightarrow\infty$, we have $p({\mathbf{y}})=\frac{1}{\pi}\delta(y_{1}^{2}+y_{2}^{2}-1)$. Since $p(R)\propto\exp^{-E^{\text{AM}}(R)}$, the score can be described as
$\begin{split}-\nabla_{R}\log p(R)&=\nabla_{R}E^{\text{AM}}(R)\\ &=2R-\frac{2\beta I_{1}(2\beta R)}{\beta I_{0}(2\beta R)}\\ &=2R-2\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)},\end{split}$ (17)
where $E^{\text{AM}}(R)$ is Eq. (15), and $I_{1}$ and $I_{0}$ are the first-order and zero-order modified Bessel functions of the first kind, respectively. With further expansion, the score becomes
$\nabla_{R}\log p(R)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\quad\Rightarrow\quad\nabla_{x}\log p(x)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\frac{x}{R}.$ (18)
We implemented Eq. (15) on a $20\times 20$ grid to obtain the exact energy and score of the model, overlaying 1000 samples from the training set. For visualization purposes, we set $\beta=20$, ensuring a well-defined energy landscape. Additionally, we normalized the energy by subtracting its minimum value to ensure that the lowest energy point remains at zero.
To align with the theoretical framework, we employed the variance-exploding (VE) SDE of the form:
@@ A.2  Experimental Details
where the diffusion coefficient $g(t)=\sigma$. Assuming $t\in(0,1]$, the variance of the diffusion kernel is given by $\int_{0}^{t}g^{2}(s)\,\mathrm{d}s=\sigma^{2}t$, enabling us to construct the corresponding Gaussian kernel:
$p({\mathbf{x}}_{t}|{\mathbf{y}})=\mathcal{N}({\mathbf{x}}_{t};{\mathbf{y}},\sigma^{2}tI)$
@@
which shares the same marginal distributions as the SDE (20) and is useful for likelihood estimation.
Empirical distribution. We represented the data as an empirical distribution
@@
where $N$ is the dimensionality of the data point $\mathbf{y}$}. Since ${\mathbf{x}}_{t}$ is drawn from the forward process distribution $p({\mathbf{x}}_{t}|{\mathbf{y}})$ that is conditioned on the data point $\mathbf{y}$}, which can be expressed as a Gaussian kernel $\mathcal{N}({\mathbf{x}}_{t};\mathbf{y},\sigma^{2}tI)$. We obtained the distribution $p(\mathbf{x}_{t},t)$ as
$\begin{split}p(\mathbf{x}_{t},t)&=\int p(\mathbf{x}_{t}\mid\mathbf{y})\,p(\mathbf{y})\,\mathrm{d}\mathbf{y}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\int\mathcal{N}\!\bigl(\mathbf{x}_{t};\mathbf{y},\sigma^{2}tI\bigr)\,\delta^{(N)}\!\bigl(\mathbf{y}-\bm{\xi}^{\mu}\bigr)\,\mathrm{d}\mathbf{y}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\mathcal{N}\!\bigl(\mathbf{x}_{t};\,\bm{\xi}^{\mu},\,\sigma^{2}t\,I\bigr)\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\frac{1}{(2\pi\,\sigma^{2}t)^{\frac{N}{2}}}\exp\!\Bigl(-\,\frac{\|\mathbf{x}_{t}-\bm{\xi}^{\mu}\|_{2}^{2}}{2\,\sigma^{2}t}\Bigr)\end{split}$ (23)
which is the derived $E^{\text{DM}}$ in the main text, when the variance of the Gaussian kernel is $\sigma^{2}t$.
Energy Computation. Following [62] we computed the log-likelihood $\log p_{\theta}({\mathbf{x}}_{0})$ given by a diffusion model, with the instantaneous change of variable formula [13], where ${\mathbf{x}}_{0}$ are the generated samples from the model. Replacing Eq. (20) into the log-likelihood equation,
$\log p_{0}({\mathbf{x}}_{0};\mathbf{\theta})=\log p_{T}({\mathbf{x}}_{T})+\int_{0}^{T}\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t$ (24)
with $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)=-\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$. The function $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ comes from the above Probability Flow ODE (21), and $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ denotes the divergence of the function (or the trace of its Jacobian).
To estimate the likelihood of the model, we integrated $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$, from a small time ${\epsilon}=10^{-5}$ to $T=1$, using a numerical integrator and added the prior logarithmic likelihood to it following Eq. (24). The divergence term $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ is computed using the Laplacian, which is computationally feasible for this toy example, instead of using the Hutchinson trace estimator [28] done in [62]. We use the RK45 method [18] implemented in the ODE solver scipy.integrate.solve_ivp from Scipy [68] to solve the above integral.
To compute the energy, we used the equation of $E^{\text{DM}}$ in the main text, which can be also computed using Eq. (24). From the Boltzmann distribution $p({\mathbf{x}})=\frac{1}{Z}\exp\bigl[-\beta E(\mathbf{x})\bigl]$, we obtained $\log p({\mathbf{x}})\propto-\beta E(\mathbf{x})$. In our case, the inverse temperature is $\beta=\frac{1}{2\sigma^{2}t}$. The equation for our VE-SDE thus becomes